# Search for gravitational waves emitted from SN 2023ixf

### v1

@@ Abstract
We present the results of a search for gravitational-wave transients associated with core-collapse supernova SN 2023ixf, which was observed in the galaxy Messier 101 via optical emission on 2023 May 19th, during the LIGO-Virgo-KAGRA 15th Engineering Run. We define a five-day on-source window during which an accompanying gravitational-wave signal may have occurred. No gravitational waves have been identified in data when at least two gravitational-wave observatories were operating, which covered 14 % of this five-day window. We report the search detection efficiency for various possible gravitational-wave emission models. Considering the distance to M101 (6.7 Mpc), we derive constraints on the gravitational-wave emission mechanism of core-collapse supernovae across a broad frequency spectrum, ranging from 50 Hz to 2 kHz where we assume the GW emission occurred when coincident data are available in the on-source window. Considering an ellipsoid model for a rotating proto-neutron star, our search is sensitive to gravitational-wave energy and luminosity for a source emitting at 50 Hz. These constraints are around an order of magnitude more stringent than those obtained so far with gravitational-wave data. The constraint on the ellipticity of the proto-neutron star that is formed is as low as 1.04, at frequencies above 1200 Hz, surpassing results from SN 2019ejj.
@@ 1 Introduction
The GW emission from CCSNe is weaker than the emission from compact binary mergers, making it detectable by the advanced generation of detectors only for nearby supernovae (Gossan et al., 2016; Szczepańczyk et al., 2021; Abbott et al., 2021c). The most likely opportunity for observations are Galactic CCSNe, but the expected rate is of the order of one or two per century (Bergh & Tammann, 1991; Cappellaro et al., 1993; Tammann et al., 1994; Diehl et al., 2006; Li et al., 2011; Adams et al., 2013). However, due to the large uncertainties of the progenitors and GW emission models, we carry out searches for GW emission from CCSNe out to distances of 20 Mpc (Abbott et al., 2016b, 2020b; Szczepańczyk et al., 2024).
SN 2023ixf was identified in Messier 101 (M101) during its rise, making it one of the closest type II CCSNe observed. The two LIGO observatories were in observing mode during the fifteenth Engineering Run (ER15) of the LIGO-Virgo-KAGRA network (Aasi et al., 2015; Acernese et al., 2015; Akutsu et al., 2021). In this letter, we report the results of the search for GWs and the new constraints on GW emission obtained with SN 2023ixf.
@@ 2.1 Summary of SN 2023ixf multi-messenger observations
SN 2023ixf (RA = 14:03:38.562, DEC = +54:18:41.94, J2000) was discovered on 2023 May 19 by Itagaki (2023) with a clear (unfiltered) magnitude of 14.9 in the host galaxy M101 (NGC 5457, Pinwheel Galaxy). M101 is at a distance of about 6.7 Mpc (see Sec. 2.3), making SN 2023ixf one of the nearest CCSNe observed in recent years. In addition, this galaxy is a well-observed object with an extensive set of pre-discovery observations. SN 2023ixf was quickly classified as a type II supernova a few hours after the discovery (Perley et al., 2023). Due to the prompt discovery and the close distance, SN 2023ixf was the target of extensive electromagnetic coverage. The optical light curve shows a rise to a maximum at about five days, followed by a plateau lasting for about one month, and a slow decline later (Hiramatsu et al., 2023; Hosseinzadeh et al., 2023; Li et al., 2024; Sgro et al., 2023; Teja et al., 2023; Yamanaka et al., 2023). The early spectroscopic observations show flash ionization features of hydrogen, helium, nitrogen, carbon and a temperature increase not explained by pure shock cooling, suggesting a delayed shock breakout in a dense circumstellar medium (Berger et al., 2023; Bersten, M. C. et al., 2024; Bostroem et al., 2023; Grefenstette et al., 2023; Chandra et al., 2024; Guetta et al., 2023; Hiramatsu et al., 2023; Hosseinzadeh et al., 2023; Koenig, 2023; Jacobson-Galan et al., 2023; Kilpatrick et al., 2023; Li et al., 2024; Martinez, L. et al., 2024; Murase, 2024; Niu et al., 2023; Pledger & Shara, 2023; Qin et al., 2024; Smith et al., 2023; Teja et al., 2023; Van Dyk et al., 2023; Vasylyev et al., 2023; Xiang et al., 2024; Yamanaka et al., 2023; Zimmerman et al., 2024). The earliest detections of X-ray and radio emission occurred four days (Grefenstette et al., 2023) and one month (Matthews et al., 2023) after the discovery, respectively. The hard X-ray (Grefenstette et al., 2023) and soft X-ray (Chandra et al., 2024) observations suggest a high and decreasing neutral hydrogen column density close to SN 2023ixf. SN 2023ixf was not detected in gamma-rays (Marti-Devesa, 2023) or in neutrinos (Thwaites et al., 2023; Nakahata & Super-Kamiokande Collaboration, 2023; Abbasi et al., 2023).
@@ 3.1 Coherent WaveBurst
We use coherent WaveBurst (cWB), a model-agnostic search algorithm, for the detection and reconstruction of transient GW signals (Klimenko et al., 2016). The algorithm identifies GW transients by searching for excess power in spectrograms and reconstructs coherent signals in multiple detectors. In previous CCSN searches (Szczepańczyk et al., 2023), spectrograms were obtained with the Wilson-Daubechies-Meyer wavelet transform (Necula et al., 2012). The SN 2023ixf analysis uses the high-resolution wavescan transform (Klimenko, 2022) that utilizes both the excess-power and cross-power statistics for the identification of GW signals and enables more accurate reconstruction of the signal waveforms. The signal detection statistic is defined as where is the total coherent energy across the detector network (Klimenko et al., 2016). To further separate GW signals from the noise, the triggers are then re-ranked with a reduced statistic , where is the XGBoost classification penalty factor which ranges between 0 (noise-like) and 1 (signal-like) (Szczepańczyk et al., 2023; Mishra et al., 2021). For this search, the XGBoost algorithm uses the SN 2023ixf sky location to improve detection sensitivity.
@@ 3.2 CCSN models
and are the moment of inertia and ellipticity of the ellipsoid, is twice the rotation frequency, is the source distance and is the inclination angle of the axis with respect to the line of sight. is a free parameter. Throughout the paper, we consider the canonical value for neutron stars (Paschalidis & Stergioulas, 2017) and we fix to to estimate the sensitivity of the search. This order-of-magnitude estimate of has been derived from simulations in which bar-mode instabilities at low rotational kinetic energy over gravitational potential energy ratio () are present (Shibagaki et al., 2020; Bugli et al., 2020).
@@ 4.1 Search result and background estimation
4.2 Detection efficiency vs distance
To evaluate the search sensitivity, we take the signal models described in Sec. 3.2 and randomize the source orientation such that it is uniformly distributed over a sphere. Then we add waveforms to the detector coincident data within the on-source window for the sky location of SN 2023ixf. We compute the search detection efficiency, defined as the fraction of detected signals with FAR lower than that of the lowest-FAR trigger found in the on-source window. Fig. 2 shows the detection efficiencies as a function of distance to a source in the direction of SN 2023ixf. We also plot the distances to the Galactic Center ( 8.5 kpc), the Large Magellanic Cloud ( 49.6 kpc) that hosted SN 1987A and the distance to M101. Table 1 shows the distances corresponding to a 50% detection efficiency for all 14 CCSN models. The distances reach up to 16.8 kpc for the non-rotating explosions. Detection capabilities for the Kur+16 s15, Pow+18 s18, Rad+19 s25, and Pan+21 NR models exceed the Galactic Center. The distances for the more extreme models are around an order of magnitude larger than for the non-rotating explosions. The Pow+23 B12 model reaches the distance of the Large Magellanic Cloud with 51.0 kpc. On the other hand, an explosion driven by a first-order quantum-chromodynamics phase transition (Kur+22 s50) an explosion driven by a first-order quantum-chromodynamics phase transition (Kur+22 s50)could be observed up to around 11.6 kpc.
@@ 4.2 Detection efficiency vs distance
Non-rotating models | Kur+16 s15 | 16.8 |
Mez+23 D15 | 5.0 | |
Oco+18 m20p | 2.0 | |
Pow+19 s18 | 9.3 | |
Pow+19 he3.5 | 4.7 | |
Rad+19 s13 | 1.2 | |
Rad+19 s25 | 10.5 | |
Pan+21 NR | 12.4 | |
Rotating models | And+19 s15fr | 3.8 |
Obe+20 Signal_O | 28.4 | |
Pan+21 SR | 38.3 | |
Pow+20 m39 | 46.7 | |
Pow+23 B12 | 51.0 | |
Phase | Kur+22 s50 | 11.6 |
@@
model |
The right panel in Fig. 2 shows the detection efficiency for long-lasting bar-mode waveforms with frequencies between 82 Hz and 2 kHz and signal durations between 1 ms and 1 s. The sensitivity distance, defined as the distance for which an efficiency of 50% is reached, increases with the signals’ peak frequency and duration, up to a few Mpc. For example, a source emitting at 2 kHz for 1 s could have been detected in this search above the threshold with an efficiency of 50% up to 1.5 Mpc. This distance reach, lower than the distance to SN 2023ixf, depends linearly on which is fixed to but could be actually larger.
@@ 5 Constraints
Assuming the GW emission occurred when coincident data are available in the on-source window, we establish constraints on several quantities characterizing a core collapse, including emitted GW energy, luminosity, and PNS ellipticity, considering the long-lasting bar-mode models.
@@ 5.1 Constraints on GW energy and luminosity
where is the source root-sum-squared GW strain for an optimally oriented source.
The GW luminosity is the ratio between the emitted GW energy and the duration of the emission. We define the duration as the time interval that contains 90% of the energy such that the GW power is given by
For the sine-Gaussians of Eq. (1) . Considering 50% detection efficiency we derive constraints on and shown in Fig. 3. The shaded region contains results from all long-lasting bar-mode models that are compared to the typical CCSN explosion energy of around erg, which can be as high as erg for hypernovae (Nomoto et al., 2010; Tanaka et al., 2009; Utrobin & Chugai, 2011). At 50 Hz the more stringent energy constraints are . Fig. 3 also shows the constraints derived from SN 2019ejj, located at 15.7 Mpc (Szczepańczyk et al., 2024) and from SN 2017eaw located at 6.7 Mpc (Abbott et al., 2020a). The constraints with SN 2023ixf are 21 times more stringent than for SN 2019ejj over the whole frequency range. For the emitted GW luminosity shown in the bottom panel, the constraints are for signals at 50 Hz and 1 s long. They are a factor of 8 more stringent compared to the SN 2019ejj over the whole frequency range.
@@ 5.2 Constraints on PNS ellipticity
Fig. 4 reports the ellipticity for a range of bar-mode GW signal frequencies and durations for a detection efficiency of 50%. The most stringent constraints on ellipticity are obtained for the signals with s, ranging from at the lowest search frequency to 1.04 at 2 kHz. The constraints get stricker with shorter signals. Over the whole frequency range, the constraints given by SN 2023ixf on the ellipticity are 2.6 more stringent than for SN 2019ejj.
@@ 6 Summary and Discussion
We present the results of a search for GW signals coincident with SN 2023ixf, which was observed during the LIGO-Virgo-KAGRA Engineering Run 15, 2023 April 24 to 2023 May 24. No significant GW candidates were identified within the 14 % of the on-source window where coincident good quality GW data are available. With different CCSN waveform models, we quantify the search sensitivity by estimating the distances at which 50% of the GW simulated signals are detected. The reported distances are up to 16.8 kpc for non-rotating explosions, and up to 51.0 kpc for rapidly rotating models. These distance sensitivities have been obtained using the FAR of 2.04 per day from the most significant event found by the search. We derive constraints on the GW energy, luminosity, and PNS ellipticity, which are the most stringent that GW detector data have achieved to date. Assuming the PNS is well modelled as a rotating triaxial ellipsoid whose moment of inertia along the rotation axis is fixed to = , we find that the ellipticity should be lower than 1.04. This value, obtained for an hypothetical 1 s-long signal at 2 kHz, is within the order of magnitude of plausible estimates derived from simulations where bar-mode instabilities are present (Obergaulinger & Aloy, 2021; Bugli et al., 2023).

### v2

@@ Abstract
We present the results of a search for gravitational-wave transients associated with core-collapse supernova SN 2023ixf, which was observed in the galaxy Messier 101 via optical emission on 2023 May 19th, during the LIGO-Virgo-KAGRA 15th Engineering Run. We define a five-day on-source window during which an accompanying gravitational-wave signal may have occurred. No gravitational waves have been identified in data when at least two gravitational-wave observatories were operating, which covered 14 % of this five-day window. We report the search detection efficiency for various possible gravitational-wave emission models. Considering the distance to M101 (6.7 Mpc), we derive constraints on the gravitational-wave emission mechanism of core-collapse supernovae across a broad frequency spectrum, ranging from 50 Hz to 2 kHz where we assume the gravitational-wave emission occurred when coincident data are available in the on-source window. Considering an ellipsoid model for a rotating proto-neutron star, our search is sensitive to gravitational-wave energy and luminosity for a source emitting at 82 Hz. These constraints are around an order of magnitude more stringent than those obtained so far with gravitational-wave data. The constraint on the ellipticity of the proto-neutron star that is formed is as low as 1.08, at frequencies above 1200 Hz, surpassing past results.
@@ 1 Introduction
The GW emission from CCSNe is weaker than the emission from compact binary mergers, making it detectable by the advanced generation of detectors only for nearby supernovae (Gossan et al., 2016; Szczepańczyk et al., 2021; Abbott et al., 2021c). The most likely opportunity for observations are Galactic CCSNe, but the expected rate is of the order of one or two per century (Bergh & Tammann, 1991; Cappellaro et al., 1993; Tammann et al., 1994; Diehl et al., 2006; Li et al., 2011; Adams et al., 2013). However, due to the large uncertainties of the progenitors and GW emission models, we carry out searches for GW emission from CCSNe out to distances of 20 Mpc (Abbott et al., 2016b, 2020; Szczepańczyk et al., 2024).
SN 2023ixf was identified in Messier 101 (M101) during its rise, making it one of the closest type II CCSNe observed. The two LIGO observatories were in observing mode during the fifteenth Engineering Run (ER15) of the LIGO-Virgo-KAGRA network (Aasi et al., 2015; Acernese et al., 2015; Akutsu et al., 2021). In this article, we report the results of the search for GWs and the new constraints on GW emission obtained with SN 2023ixf.
@@ 2.1 Summary of SN 2023ixf multi-messenger observations
SN 2023ixf (RA = 14:03:38.562, DEC = +54:18:41.94, J2000) was discovered on 2023 May 19 by Itagaki (2023) with a clear (unfiltered) magnitude of 14.9 in the host galaxy M101 (NGC 5457, Pinwheel Galaxy). M101 is at a distance of about 6.7 Mpc (see Sec. 2.3), making SN 2023ixf one of the nearest CCSNe observed in recent years. In addition, this galaxy is a well-observed object with an extensive set of pre-discovery observations. SN 2023ixf was quickly classified as a type II supernova a few hours after the discovery (Perley et al., 2023). Due to the prompt discovery and the close distance, SN 2023ixf was the target of extensive electromagnetic coverage. The optical light curve shows a rise to a maximum at about five days, followed by a plateau lasting for about one month, and a slow decline later (Hiramatsu et al., 2023; Hosseinzadeh et al., 2023; Li et al., 2024; Sgro et al., 2023; Teja et al., 2023; Yamanaka et al., 2023). The early spectroscopic observations show flash ionization features of hydrogen, helium, nitrogen, carbon and a temperature increase not explained by pure shock cooling, suggesting a delayed shock breakout in a dense circumstellar medium (Berger et al., 2023; Bersten, M. C. et al., 2024; Bostroem et al., 2023; Grefenstette et al., 2023; Chandra et al., 2024; Guetta et al., 2023; Hiramatsu et al., 2023; Hosseinzadeh et al., 2023; Koenig, 2023; Jacobson-Galan et al., 2023; Kilpatrick et al., 2023; Li et al., 2024; Martinez, L. et al., 2024; Murase, 2024; Niu et al., 2023; Pledger & Shara, 2023; Qin et al., 2024; Smith et al., 2023; Teja et al., 2023; Van Dyk et al., 2023; Vasylyev et al., 2023; Xiang et al., 2024; Yamanaka et al., 2023; Zimmerman et al., 2024). The earliest detections of X-ray and radio emission occurred four days (Grefenstette et al., 2023) and one month (Matthews et al., 2023) after the discovery, respectively. The hard X-ray (Grefenstette et al., 2023) and soft X-ray (Chandra et al., 2024; Panjkov et al., 2024) observations suggest a high and decreasing neutral hydrogen column density close to SN 2023ixf. SN 2023ixf was not detected in gamma-rays (Marti-Devesa, 2023) or in neutrinos (Thwaites et al., 2023; Nakahata & Super-Kamiokande Collaboration, 2023; Abbasi et al., 2023).
@@ 3.1 Coherent WaveBurst
We use coherent WaveBurst (cWB), a model-agnostic search algorithm, for the detection and reconstruction of transient GW signals (Klimenko et al., 2016). The algorithm identifies GW transients by searching for excess power in spectrograms and reconstructs coherent signals in multiple detectors. In previous CCSN searches (Szczepańczyk et al., 2023), spectrograms were obtained with the Wilson-Daubechies-Meyer wavelet transform (Necula et al., 2012). The SN 2023ixf analysis uses the high-resolution wavescan transform (Klimenko, 2022) that utilizes both the excess-power and cross-power statistics for the identification of GW signals and enables more accurate reconstruction of the signal waveforms Mishra et al. (2025). The signal detection statistic is defined as where is the total coherent energy across the detector network (Klimenko et al., 2016). To further separate GW signals from the noise, the triggers are then re-ranked with a reduced statistic , where is the XGBoost classification penalty factor which ranges between 0 (noise-like) and 1 (signal-like) (Szczepańczyk et al., 2023; Mishra et al., 2021). For this search, the XGBoost algorithm uses the SN 2023ixf sky location to improve detection sensitivity.
@@ 3.2 CCSN models
and are the moment of inertia and ellipticity of the ellipsoid, is twice the rotation frequency, is the source distance and is the inclination angle of the axis with respect to the line of sight. is a free parameter. Throughout the paper, we consider the canonical value for neutron stars (Paschalidis & Stergioulas, 2017) and keep as a free parameter.
@@ 4.1 Search result and background estimation
4.2 Detection efficiency
To evaluate the search sensitivity, we take the signal models described in Sec. 3.2 and randomize the source orientation such that it is uniformly distributed over a sphere. Then we add waveforms to the detector coincident data within the on-source window for the sky location of SN 2023ixf. We compute the search detection efficiency, defined as the fraction of detected signals with FAR lower than event in years. This FAR corresponds to a false alarm probability of . At the distance of SN 2023ixf none of the 14 models from numerical simulations are detected. In Table 1 we report the distance at which we recover 90% of the added signals for all 14 CCSN models. The distances reach up to 6.9 kpc for the non-rotating explosions, which means that signals from non-rotating or slow-rotating progenitor CCSN could be missed within the Galaxy for sources further than the Galactic Center. The distances for the more extreme models are around a factor 4 larger than for the non-rotating explosions, exceeding the Galaxy boundaries – 29.9 kpc for Pow+23 B12 model – but without reaching the large Magellanic Cloud distance. Finally, for the explosion driven by a first-order quantum-chromodynamics phase transition, a detection of 90% is never achieved because only one polarization is extracted from the 2D numerical simulations.
@@ 4.2 Detection efficiency
| Non-rotating models | Kur+16 s15 | 6.9 |
| Mez+23 D15 | 2.9 | |
| Oco+18 m20p | 1.0 | |
| Pow+19 s18 | 5.5 | |
| Pow+19 he3.5 | 2.8 | |
| Rad+19 s13 | 0.6 | |
| Rad+19 s25 | 5.8 | |
| Pan+21 NR | 6.6 | |
| Rotating models | And+19 s15fr | 1.8 |
| Obe+20 Signal_O | 13.4 | |
| Pan+21 SR | 18.2 | |
| Pow+20 m39 | 19.6 | |
| Pow+23 B12 | 29.9 | |
| Phase | Kur+22 s50 | |
@@
| model | ||
Fig. 2 shows the detection efficiency for long-lasting bar-mode waveforms with frequencies between 82 Hz and 2 kHz and signal durations between 1 ms and 1 s. The right plot is as a function of for a source at the location of SN 2023ixf. The left plot is as a function of distance for . The sensitivity increases with the signals peak frequency and duration. For instance, we could detect at 90% confidence level a signal lasting 1 s at 2 kHz for . For lower frequencies, if we assume the canonical value , the source would need to be highly deformed ().
@@ 5 Constraints
Assuming the GW emission occurred when coincident data are available in the on-source window, we establish constraints on several quantities characterizing a core collapse, including emitted GW energy, luminosity, and PNS ellipticity, considering the long-lasting bar-mode model. These bar-mode instabilities are sometimes present in simulations at low rotational kinetic energy over gravitational potential energy ratio ().
@@ 5.1 Constraints on GW energy and luminosity
where the GW strain squared integral is computed for an optimally oriented source.
The GW luminosity is the ratio between the emitted GW energy and the duration of the emission. We define the duration as the time interval that contains 90% of the energy such that the GW average luminosity is given by
For the sine-Gaussians of Eq. (1) . Considering the value corresponding to 90% detection efficiency we derive constraints on and shown in Fig. 3. The shaded region contains results from all long-lasting bar-mode models. At 82 Hz the more stringent energy constraints are . Fig. 3 also shows the constraints derived from GW searches targetting CCSNe during the third observing run (O3) of the LIGO-Virgo-KAGRA network. The constraints with SN 2023ixf are 49 times more stringent than for O3 CCSNe over the whole frequency range. For the emitted GW average luminosity shown in the bottom panel, the constraints are for signals at 82 Hz and 1 s long. They are a factor of 36 more stringent compared to the O3 CCSNe over the whole frequency range. This upper limit is between 2 and 5 orders of magnitude larger than the GW average luminosity predicted by numerical simulations for the different explosion mechanisms (Szczepańczyk et al., 2021) showing that we could soon exclude some of the most optimistically luminous models with a closer CCSN. It is also orders of magnitude larger than the bolometric luminosity reported in Zimmerman et al. (2024).
@@ 5.2 Constraints on PNS ellipticity
Fig. 4 reports the ellipticity for a range of bar-mode GW signal frequencies and durations for a detection efficiency of 90%. The most stringent constraints on ellipticity are obtained for the signals with s, ranging from at the lowest search frequency to 1.08 at 2 kHz. The constraints get stricter with longer signals. Over the whole frequency range, the constraints given by SN 2023ixf on the ellipticity are 6.8 more stringent than for O3 CCSNe.
@@ 6 Summary and Discussion
We present the results of a search for GW signals coincident with SN 2023ixf, which was observed during the LIGO-Virgo-KAGRA Engineering Run 15, 2023 April 24 to 2023 May 24. No significant GW candidates were identified within the 14 % of the on-source window where coincident good quality GW data are available. With different CCSN waveform models, we quantify the search sensitivity by estimating the distances at which 90% of the GW simulated signals are detected. The reported distances are up to 6.9 kpc for non-rotating explosions, and up to 29.9 kpc for rapidly rotating models. These distance sensitivities have been obtained using the FAR of event in years . We derive constraints on the GW energy, luminosity, and PNS ellipticity, which are the most stringent that GW detector data have achieved to date. Assuming the PNS is well modelled as a rotating triaxial ellipsoid whose moment of inertia along the rotation axis is fixed to = , we find that the ellipticity should be lower than 1.08. This value, obtained for an hypothetical 1 s-long signal at 2 kHz, is one order of magnitude larger than plausible estimates () derived from simulations where bar-mode instabilities are present (Shibagaki et al., 2021; Obergaulinger & Aloy, 2021; Bugli et al., 2023).